7. The probability that a man aged 50 will survive one year is 98428. Show that the probability that, out of 5 men aged 50, 3 at least will die within a year is 0000385
Answer #1

P(atleast 3 ) = P(X=3)+P(X=4)+P(X=5)

P(X=3) = C(5,3)*(0.98428)^2*(1-0.98428)^3 =

0.00003763526

P(X=4) = C(5,4)*(0.98428)*(1-0.98428)^4 =

0.00000030054

P(X=5) = C(5,5)*(1-0.98428)^5 =

0.00000000096

P(X>=3) = 0.0000385